# On the continuous properties for the 3D incompressible rotating Euler equations

### v1

@@ 1.3 Main Result
Furthermore, if $r<\infty$, the above solution $u\in C([0,T];B^{s}_{p,r})$ is continuous dependence in $B^{s}_{p,r}$.
@@
Next, we consider the the property of continuous dependence of solutions of the Cauchy problem for (1.2) and show that the solution map for the Euler-Coriolis system (1.2) cannot be uniformly continuous in Besov spaces $B^{s}_{p,r}$. From now on, we denote
@@
Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying (1.3). The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) is not uniformly continuous from any bounded subset in $B^{s}_{p,r}$ into $\mathcal{C}([0,T];B^{s}_{p,r})$. More precisely, there exist two initial data sequences $f_{n}+g_{n}$ and $f_{n}$ such that
$\left\|(f_{n})^{(2)}\right\|_{B^{s}_{p,r}}\approx 1\quad\text{and}\quad\lim_{n\rightarrow\infty}\left\|(f_{n})^{(1)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\left\|(f_{n})^{(3)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\|g_{n}\|_{B^{s}_{p,r}}=0$
@@
We should mention that Theorem 1.2 is stronger than that for the Euler equations in [19]. On the one hand, the non-uniform continuous is established in the more framework of Besov spaces. On the other hand, Theorem 1.2 is focused on the second component of data-to-solution map which leads to the non-uniform continuous.
@@
Assume that $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ with $(s,p,r)$ satisfying (1.3). For any $\alpha\in(0,1)$, there exits $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.2) satisfying
@@
Assume that $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ with $(s,p,r=\infty)$ satisfying (1.3). There exits $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.2) satisfies
@@
As mentioned above, (1.2) with $\Omega=0$ is the classical Euler equations. Thus Our main Theorems 1.1-1.4 holds true for the classical Euler equations.
@@ 1.4 Strategies to the proof
Proof of Theorem 1.2. The main task is to construct initial data such that the second component of the initial velocity is uniformly bounded in $B^{s}_{p,r}$, and in the same time to have the other component of the initial velocity remain small. We first choose two sequences of initial values $u^{n}_{0,1}=f_{n}+g_{n}$ and $u^{n}_{0,2}=f_{n}+g_{n}$, which can generate the approximate solutions $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$, respectively. Next, we perform perturbation analysis (See Proposition 3.1) to show the second component of the difference between the data-to-solution maps $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$ is bounded below by a positive constant at any later time, which means the solution map is not uniformly continuous.
Proof of Theorem 1.3. In order to prove the failure of Hölder regularity of solutions for the (1.2), we would like to first find initial condition to generate the corresponding solutions to (1.2) satisfying that some positive constant $C$
$\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}\geq Cn.$
Our key idea is to decompose the difference between the data-to-solution maps $\mathbf{S}_{t}(u_{0})$ and $\mathbf{S}_{t}(u_{0})$ as
$\mathbf{S}_{t}(u_{0})-u_{0}=\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})-t\mathbf{v}_{0}(u_{0})\quad\text{and}\quad\mathbf{v}_{0}(u_{0})=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0}).$
There remains two main difficulties that we need to overcome.
@@
To carry out the analysis for these above two points, a big part of the technical difficulties lies in the construction of our initial data. We construct a new initial data which is somewhat different from Theorem 1.2. Finally, by utilizing the commutator estimate and some basic analysis, we hope to obtain the the loss of Hölder regularity of solutions to (1.2).
@@ 1.5 Organization of our paper
  • In Section 3, we show the non-uniform continuous of the solution map (Theorem 1.2).
@@
  • In Section 5, we establish the discontinuous at zero time of the solution map (Theorem 1.4).
@@ 2 Preliminaries
    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-d}\int_{\mathbb{R}^{3}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
  • We denote the projection
@@
    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant 3}$, where $\mathcal{P}_{ij}\equiv\delta_{ij}+R_{i}R_{j}$ with $\delta_{ij}$ being the Kronecker delta ($\delta_{ij}=0$ for $i\neq j$ and $\delta_{ii}=0$) and $R_{i}$ being the Riesz transform with symbol $-\mathrm{i}\xi_{1}/|\xi|$. Obviously, $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$, and if $\mathord{{\rm div}}\,u=\mathord{{\rm div}}\,v=0$, it holds that $\mathcal{Q}(u\cdot\nabla v)=\mathcal{Q}(v\cdot\nabla u).$
@@
$\forall\,u\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{3}),\quad\Delta_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
@@
Definition 2.1 ([3]).
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the nonhomogeneous and homogeneous Besov spaces, respectively,
$B^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}:\;\|f\|_{B^{s}_{p,r}}:=\left\|2^{js}\|\Delta_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\geq-1)}<\infty\right\}$
and
@@
Assume that $(s,p,r)$ satisfies (1.3). Then there exists a constant $C$, depending only on $d,p,r,s$, such that for all $u,v\in B^{s}_{p,r}$ with $\mathrm{div\,}u=\mathrm{div\,}v=0$,
$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\|u\|_{B^{s}_{p,r}}\|v\|_{B^{s}_{p,r}}$
and
@@
We assume that $\widehat{\phi}$ is even and real-valued such that $\theta$ is a real-valued function. Then we can define the new real-valued function
@@
$\mathrm{supp}\ \widehat{\theta}\subset\left\{\xi=(\xi_{1},\xi_{2},\xi_{3})\in\mathbb{R}^{3}:\ |\xi|\leq\frac{1}{2}\right\}.$ (3.6)
@@
$\mathrm{supp}\ \mathcal{F}\Big{(}\theta(x)T(\lambda x_{1})\Big{)}\subset\left\{\xi\in\mathbb{R}^{3}:\ \lambda-\frac{1}{2}\leq|\xi|\leq\lambda+\frac{1}{2},\ |\xi_{2}|\leq\frac{1}{2},\ |\xi_{3}|\leq\frac{1}{2}\right\}.$ (3.7)
@@
From (3.7) and Bernstein’s inequality (see Lemma 2.1), it follows that the desired results of Lemma 3.1. ∎
@@
The definitions of $f_{n}$ and $g_{n}$ tell directly us that
$2^{ns}g_{n}\cdot\nabla f_{n}^{(2)}$ $=\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
$\quad+2^{-n}\phi(x_{1})\phi^{\prime}(x_{1})\sin\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
$\quad-2^{-2n}\phi(x_{1})\phi^{\prime\prime}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})$
$\quad+2^{-2n}\left[\phi^{\prime}(x_{1})\right]^{2}\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3}).$
Then, we have
@@
here and in what follows we denote $\mathbf{v}_{0}(u_{0}):=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0})$ and
@@
$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s-1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s-1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\leq C\int^{t}_{0}\left(\|u\|_{B^{s}_{p,r}}\|u\|_{B^{s-1}_{p,r}}+\|u\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\leq Ct\|u_{0}\|_{B^{s-1}_{p,r}},$ (3.13)
@@
$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s}_{p,r}}\right)\mathrm{d}\tau$
$\leq Ct\left(1+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+1}_{p,r}}\right),$ (3.14)
@@
$\leq\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s+1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s+1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s+1}_{p,r}}\right)\mathrm{d}\tau$
$\leq Ct\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$ (3.15)
@@
$\leq\int^{t}_{0}\|\mathcal{Q}(u)-\mathcal{Q}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\cdot\nabla u-u_{0}\cdot\nabla u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\quad+\Omega\int^{t}_{0}\|\mathcal{P}\left(e_{3}\times(u-u_{0})\right)\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\|u(\tau)\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
$\quad\ +\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\|u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\|u_{0}\|_{B^{s+1}_{p,r}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
$\quad\ +\|u_{0}\|_{L^{\infty}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\mathrm{d}\tau.$ (3.16)
@@
$\mathbf{S}_{t}(u^{n}_{0})$ $=\mathbf{S}_{t}(u^{n}_{0})-u^{n}_{0}+t\mathbf{v}_{0}(u^{n}_{0})-(f_{n}+g_{n})-t(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})$
@@
$\mathbf{S}_{t}(f_{n})$ $=\mathbf{S}_{t}(f_{n})-f_{n}+t\mathbf{v}_{0}(f_{n})-f_{n}-tf_{n}\cdot\nabla f_{n}$
@@
$\left\|\mathbf{S}^{(2)}_{t}(u^{n}_{0})-\mathbf{S}^{(2)}_{t}(f_{n})\right\|_{B^{s}_{p,r}}$ $\geq t\left\|\left[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\right]^{(2)}\right\|_{B^{s}_{p,r}}-t\|\mathcal{P}(\Omega e_{3}\times g_{n})\|_{B^{s}_{p,r}}$
$\quad-t\left\|\mathcal{Q}\big{[}(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\big{]}\right\|_{B^{s}_{p,r}}-\|g_{n}\|_{B^{s}_{p,r}}-Ct^{2}$
$\geq t\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,r}}-Ct^{2}-C2^{-n},$ (3.17)
@@
Assume that $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. Define the divergence-free vector field $u_{0}$ as above. Then there exists some sufficiently large $n\in\mathbb{Z}^{+}$ and some positive constant $c_{0}$ such that
@@
Due to the simple fact (3.7), it holds that for some sufficiently large $n\in\mathbb{Z}^{+}$
@@
×$\quad\times\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi^{2}(x_{3})\phi(x_{2})\phi^{\prime}(x_{2})+\text{Remainder terms}.$
For the Remainder terms, noticing that at most one partial derivative falls on cosine function, thus we find easily that
@@
Thus we have for some sufficiently large $n\in\mathbb{Z}^{+}$
$n^{2}2^{n(s+1)}\left\|u_{0}^{(1)}{\bf\partial_{1}\partial_{1}}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\geq c_{0}2^{2n}-C2^{n}.$ (4.21)
@@
Noticing that (4.20), from (4.21) and (4.22), we obtain that for some sufficiently large $n\in\mathbb{Z}^{+}$
@@
$\lesssim\int^{t}_{0}\|\mathcal{P}\left(u\cdot\nabla u-u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times(u-u_{0})\right)\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\left(\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}+\|u-u_{0}\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\lesssim\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t^{2}.$

### v2

@@ 1.3 Main Result
Furthermore, if $r<\infty$, the above solution $u\in C([0,T];B^{s}_{p,r})$ is continuous dependent in $B^{s}_{p,r}$.
@@
Next, we consider the property of continuous dependence of solutions of the Cauchy problem for (1.1) and show that the solution map for the Euler-Coriolis system (1.1) cannot be uniformly continuous in Besov spaces $B^{s}_{p,r}$. From now on, we denote
@@
and denote any bounded subset $U_{R}$ in $B^{s}_{p,r}(\mathbb{R}^{3})$ by
$U_{R}:=\left\{u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3}):\|u_{0}\|_{B^{s}_{p,r}(\mathbb{R}^{3})}\leq R,\;\mathrm{div}u_{0}=0\right\}.$
@@
Assume that $(s,p,r)$ satisfying (1.3). The data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.1) is not uniformly continuous from any bounded subset in $U_{R}\subset B^{s}_{p,r}$ into $\mathcal{C}([0,T];B^{s}_{p,r})$. More precisely, there exist two initial data sequences $f_{n}+g_{n}$ and $f_{n}$ such that
$\left\|(f_{n})^{(2)}\right\|_{B^{s}_{p,r}}\approx 1\quad\text{and}$
$\lim_{n\rightarrow\infty}\left\|(f_{n})^{(1)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\left\|(f_{n})^{(3)}\right\|_{B^{s}_{p,r}}=\lim_{n\rightarrow\infty}\|g_{n}\|_{B^{s}_{p,r}}=0,$
@@
We should mention that Theorem 1.2 is stronger than that for the Euler equations in [19]. On the one hand, the non-uniform continuous is established in the more general framework of Besov spaces. On the other hand, Theorem 1.2 is focused on the second component of data-to-solution map which leads to the non-uniform continuous.
@@
Assume that $(s,p,r)$ satisfying (1.3). For any $\alpha\in(0,1)$, there exists $u_{0}\in B^{s}_{p,r}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})\in\mathcal{C}([0,T];B^{s}_{p,r})$ of the Cauchy problem (1.1) satisfies
@@
Assume that $(s,p,r=\infty)$ satisfying (1.3). There exists $u_{0}\in B^{s}_{p,\infty}(\mathbb{R}^{3})$ such that the data-to-solution map $u_{0}\mapsto\mathbf{S}_{t}(u_{0})$ of the Cauchy problem (1.1) satisfies
@@
As mentioned above, (1.1) with $\Omega=0$ is the classical Euler equations. Thus our main Theorems 1.1-1.4 hold true for the classical Euler equations.
@@ 1.4 Strategies to the proof
Proof of Theorem 1.2. The main task is to construct initial data such that the second component of the initial velocity is uniformly bounded in $B^{s}_{p,r}$, and in the same time to have the other component of the initial velocity remain small. We first choose two sequences of initial values $u^{n}_{0,1}=f_{n}+g_{n}$ and $u^{n}_{0,2}=f_{n}$, which can generate the approximate solutions $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$, respectively. Next, we perform perturbation analysis (See Proposition 3.1) to show the second component of the difference between the data-to-solution maps $\mathbf{S}_{t}(u^{n}_{0,1})$ and $\mathbf{S}_{t}(u^{n}_{0,2})$ is bounded below by a positive constant at any later time, which means the solution map is not uniformly continuous.
Proof of Theorem 1.3. In order to prove the failure of Hölder regularity of solutions for the (1.1), we would like to first find initial condition to generate the corresponding solutions to (1.1) satisfying that
$\frac{\left\|\mathbf{S}_{t}(u_{0})-u_{0}\right\|_{B^{s}_{p,r}}}{t^{\alpha}}\geq C(t)\to+\infty\quad\text{as}\quad t\to 0^{+}.$
Our key idea is to decompose the difference between the data-to-solution map $\mathbf{S}_{t}(u_{0})$ and initial data $u_{0}$ as
$\mathbf{S}_{t}(u_{0})-u_{0}=\mathbf{S}_{t}(u_{0})-u_{0}+t\mathbf{v}_{0}(u_{0})-t\mathbf{v}_{0}(u_{0}),$
here and in what follows we denote $\mathbf{v}_{0}(u_{0}):=\mathcal{P}(u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times u_{0})$.
There remains two main difficulties that we need to overcome
@@
To carry out the analysis for these above two points, a big part of the technical difficulties lies in the construction of our initial data. We construct a new initial data which is somewhat different from Theorem 1.2. Finally, thanks to commutator estimates and some basic analysis, we obtain the loss of Hölder regularity of solutions to (1.1).
@@ 1.5 Organization of our paper
  • In Section 3, we show the non-uniform continuity of the solution map (Theorem 1.2).
@@
  • In Section 5, we establish the discontinuity at zero time of the solution map (Theorem 1.4).
@@ 2 Preliminaries
    $u(x)=\mathcal{F}^{-1}\widehat{u}(x)=(2\pi)^{-3}\int_{\mathbb{R}^{3}}e^{\mathrm{i}x\cdot\xi}\widehat{u}(\xi)\mathrm{d}\xi.$
  • We denote the Leray projection
@@
    In $\mathbb{R}^{d}$, $\mathcal{P}$ can be defined by $\mathcal{P}=\mathrm{Id}+(-\Delta)^{-1}\nabla{\rm{div}}$, or equivalently, $\mathcal{P}=(\mathcal{P}_{ij})_{1\leqslant i,j\leqslant 3}$, where $\mathcal{P}_{ij}\equiv\delta_{ij}+R_{i}R_{j}$ with $\delta_{ij}$ being the Kronecker delta ($\delta_{ij}=0$ for $i\neq j$ and $\delta_{ii}=0$) and $R_{i}$ being the Riesz transform with symbol $\mathrm{i}\xi_{i}/|\xi|$. Obviously, $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$, and if $\mathord{{\rm div}}\,u=\mathord{{\rm div}}\,v=0$, it holds that $\mathcal{Q}(u\cdot\nabla v)=\mathcal{Q}(v\cdot\nabla u).$
@@
$\forall\,u\in\mathcal{S}^{\prime}_{h}(\mathbb{R}^{3}),\quad\dot{\Delta}_{j}u=\varphi(2^{-j}D)u,\;\;\text{if}\;j\in\mathbb{Z},$
@@
Definition 2.1 (Nonhomogeneous Besov spaces, see [3]).
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the nonhomogeneous Besov spaces
$B^{s}_{p,r}:=\left\{f\in\mathcal{S}^{\prime}:\;\|f\|_{B^{s}_{p,r}}:=\left\|2^{js}\|\Delta_{j}f\|_{L_{x}^{p}}\right\|_{\ell^{r}(j\geq-1)}<\infty\right\}.$
Definition 2.2 (Homogeneous Besov spaces, see [3]).
Let $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. We define the homogeneous Besov spaces
@@
Assume that $(s,p,r)$ satisfies (1.3). Then there exists a constant $C$, depending only on $d,p,r,s$, such that for all $u,v\in B^{s}_{p,r}$ with $\mathrm{div\,}u=\mathrm{div\,}v=0$
$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\left(\|u\|_{C^{0,1}}\|v\|_{B^{s}_{p,r}}+\|v\|_{C^{0,1}}\|u\|_{B^{s}_{p,r}}\right)$
with $\mathcal{Q}=-(-\Delta)^{-1}\nabla{\rm{div}}$. Furthermore, there holds
$\|\mathcal{Q}(u\cdot\nabla v)\|_{B^{s}_{p,r}}\leq C\|u\|_{B^{s}_{p,r}}\|v\|_{B^{s}_{p,r}},$
@@
The inverse Fourier transform allows us to recover $\phi$ from $\widehat{\phi}$, namely, $\phi(x)=\mathcal{F}^{-1}(\widehat{\phi}(\xi))$. Then we can define the new real-valued function
@@
[…]
@@
$\mathrm{supp}\ \mathcal{F}\Big(\theta(x)T(\lambda x_{1})\Big)\subset\left\{\xi\in\mathbb{R}^{3}:\ \lambda-\frac{1}{8}\leq|\xi_{1}|\leq\lambda+\frac{1}{8},\ |\xi_{2}|\leq\frac{1}{8},\ |\xi_{3}|\leq\frac{1}{8}\right\}.$ (3.7)
@@
Remark 3.1.
The choice of $\frac{17}{12}\in(\frac{4}{3},\frac{3}{2})$ is crucial but not unique. In fact, due to (3.7), which implies that
$\mathrm{supp}\ \mathcal{F}\left(f_{n}\right)\subset\left\{\xi\in\mathbb{R}^{3}:\ \frac{17}{12}2^{n}-1\leq|\xi|\leq\frac{17}{12}2^{n}+1\right\},$
and notice that $\varphi(2^{-j}\xi)\equiv 1$ in $\left\{\xi\in\mathbb{R}^{3}:\ \frac{4}{3}2^{j}\leq|\xi|\leq\frac{3}{2}2^{j}\right\},$ we have $\mathcal{F}(\dot{\Delta}_{j}f_{n})=\varphi(2^{-j}\cdot)\widehat{f_{n}}=0$ for $j\neq n,$ and thus $\dot{\Delta}_{j}f_{n}=f_{n}$ if $j=n$.
@@
From Remark 3.1 and Bernstein’s inequality (see Lemma 2.1), we obtain the desired results of Lemma 3.1. ∎
@@
The definitions of $f_{n}$ and $g_{n}$ directly tell us that
$2^{ns}g_{n}\cdot\nabla f_{n}^{(2)}$ $=-\left(\frac{17}{12}\right)^{2}\phi^{2}(x_{1})\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi(x_{2})\phi^{\prime}(x_{2})\phi^{2}(x_{3})+\mathbf{R}.$
For the remainder terms $\mathbf{R}$, noticing that at most one partial derivative falls on cosine function, then $\|\mathbf{R}\|_{L^{p}}\leq C2^{-n}$. Thus, we have
@@
here and in what follows we denote
@@
$\leq$ $~\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s-1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s-1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\leq$ $~C\int^{t}_{0}\left(\|u\|_{B^{s}_{p,r}}\|u\|_{B^{s-1}_{p,r}}+\|u\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\leq$ $~Ct\|u_{0}\|_{B^{s-1}_{p,r}},$ (3.13)
@@
$\leq$ $~\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s}_{p,r}}\right)\mathrm{d}\tau$
$\leq$ $~Ct\left(1+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+1}_{p,r}}\right),$ (3.14)
@@
$\leq$ $~\int^{t}_{0}\left(\|u\cdot\nabla u\|_{B^{s+1}_{p,r}}+\|\mathcal{P}(\Omega e_{3}\times u)\|_{B^{s+1}_{p,r}}+\|\mathcal{Q}(u\cdot\nabla u)\|_{B^{s+1}_{p,r}}\right)\mathrm{d}\tau$
$\leq$ $~Ct\left(\|u_{0}\|_{B^{s+1}_{p,r}}+\|u_{0}\|_{B^{s-1}_{p,r}}\|u_{0}\|_{B^{s+2}_{p,r}}\right).$ (3.15)
@@
$\leq$ $~\int^{t}_{0}\|\mathcal{Q}(u)-\mathcal{Q}(u_{0})\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u\cdot\nabla u-u_{0}\cdot\nabla u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$+\Omega\int^{t}_{0}\|\mathcal{P}\left(e_{3}\times(u-u_{0})\right)\|_{B^{s}_{p,r}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\|u(\tau)\|_{B^{s+1}_{p,r}}\mathrm{d}\tau$
$+\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\|u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s}_{p,r}}\mathrm{d}\tau+\|u_{0}\|_{B^{s+1}_{p,r}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{L^{\infty}}\mathrm{d}\tau$
$+\|u_{0}\|_{L^{\infty}}\int^{t}_{0}\|u(\tau)-u_{0}\|_{B^{s+1}_{p,r}}\mathrm{d}\tau.$ (3.16)
@@
$\mathbf{S}_{t}(u^{n}_{0})$ $=\mathbf{S}_{t}(u^{n}_{0})-u^{n}_{0}+t\mathbf{v}_{0}(u^{n}_{0})+(f_{n}+g_{n})-t(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})$
@@
$\mathbf{S}_{t}(f_{n})$ $=\mathbf{S}_{t}(f_{n})-f_{n}+t\mathbf{v}_{0}(f_{n})+f_{n}-tf_{n}\cdot\nabla f_{n}$
@@
$\left\|\mathbf{S}^{(2)}_{t}(u^{n}_{0})-\mathbf{S}^{(2)}_{t}(f_{n})\right\|_{B^{s}_{p,r}}$
$\geq$ $~t\left\|\left[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\right]^{(2)}\right\|_{B^{s}_{p,r}}-t\|\mathcal{P}(\Omega e_{3}\times g_{n})\|_{B^{s}_{p,r}}$
$-t\left\|\mathcal{Q}\big[(f_{n}+g_{n})\cdot\nabla(f_{n}+g_{n})-f_{n}\cdot\nabla f_{n}\big]\right\|_{B^{s}_{p,r}}-\|g_{n}\|_{B^{s}_{p,r}}-Ct^{2}$
$\geq$ $~t\left\|g_{n}\cdot\nabla f_{n}^{(2)}\right\|_{B^{s}_{p,r}}-Ct^{2}-C2^{-n},$ (3.17)
@@
Assume that $s\in\mathbb{R}$ and $(p,r)\in[1,\infty]^{2}$. Define the divergence-free vector field $u_{0}$ as above. Then there exists some sufficiently large $n\in\mathbb{N}$ and some positive constant $c_{0}$ such that
@@
Due to the simple fact (3.7), it holds that for some sufficiently large $n\in\mathbb{N}$
@@
$\quad\times\cos\left(\frac{17}{12}2^{n}x_{1}\right)\phi^{2}(x_{3})\phi(x_{2})\phi^{\prime}(x_{2})+\text{Remainder terms}.$
For the Remainder terms, noticing that at most one partial derivative falls on cosine function, thus we easily find that
@@
Thus we have for some sufficiently large $n\in\mathbb{N}$
$n^{2}2^{n(s+1)}\left\|u_{0}^{(1)}\partial_{1}\partial_{1}a_{n}\right\|_{L^{p}(\mathbb{R}^{3})}$ $\geq c_{0}2^{2n}-C2^{n}.$ (4.21)
@@
Noticing that (4.20), from (4.21) and (4.22), we obtain that for some sufficiently large $n\in\mathbb{N}$
@@
$\lesssim$ $~\int^{t}_{0}\|\mathcal{P}\left(u\cdot\nabla u-u_{0}\cdot\nabla u_{0}+\Omega e_{3}\times(u-u_{0})\right)\|_{{B}^{s-2}_{p,r}}\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\left(\|(u-u_{0})(u+u_{0})\|_{B^{s-1}_{p,r}}+\|u-u_{0}\|_{B^{s-1}_{p,r}}\right)\mathrm{d}\tau$
$\lesssim$ $~\int^{t}_{0}\|u-u_{0}\|_{B^{s-1}_{p,r}}\mathrm{d}\tau\lesssim t^{2}.$